Two men are 50 and 30 years old. How many years ago was the older twice the age of the younger?
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

Two men are 50 and 30 years old. How many years ago was the older twice the age of the younger?
-----
Let that number of years be "x":
----
Equation:
50+x = 2(30+x)
----
50+x = 60 + 2x
x = -10
---
Ans: 10 years ago.
================
Cheers,
Stan H.
================

Answer by solver91311(24713)



Let represent the "number of years ago". The older guy is at that time, and the younger is at that time and we know that:



Solve for

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism